# Slab tilings, flips and the triple twist

### v1

@@ 6 Final remarks
The fourth is regarding other moves besides the flip. In the domino case, there is also the trit, and the question on whether all tilings of a given region are connected by flips and trits is open. In the slab case, the existence of a set of local moves which connect all or most tilings of a given region is hindered by Proposition 5.8. In the example of Figure 21, however, it is possible to perform moves involving a quantity of pieces proportional to the lengths of the sides of the box. It can also be asked how common is a situation comparable to Proposition 5.8. It would be interesting to clarify the situation in greater generality.
@@
Figure 30: A tiling ${\mathbf{t}}$ with $\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})=\pm 1$.

### v2

@@ 8 Final remarks
The fourth is regarding other moves besides the flip. In the domino case, there is also the trit, and the question on whether all tilings of a given region are connected by flips and trits is open. In the slab case, the existence of a set of local moves which connect all or most tilings of a given region is hindered by Remark 6.5. In the example of Figure 21, however, it is possible to perform moves involving a quantity of pieces proportional to the lengths of the sides of the box. It can also be asked how common is a situation comparable to the example in Lemma 6.3 and Corollary 6.4. It would be interesting to clarify the situation in greater generality.
@@
Figure 30: A tiling ${\mathbf{t}}$ with $\operatorname{Tw}_{\kappa_{z}}({\mathbf{t}})=\pm 1$
@@
9 Declarations
9.1 Funding and Acknowledgements
The second author is thankful for the generous support of CNPq, CAPES and FAPERJ (Brazil). The first and third authors are thankful to Projeto Arquimedes and to CNPq for the support in the PIBIC undergraduate project.